# Forecasting intraday foreign exchange volatility with functional GARCH approaches

### v2

@@ 7 Conclusion
This paper seeks to predict the conditional volatility of intraday return curves in FX markets using functional GARCH-type models. Our dataset includes the intraday 5-minute frequency spot price and bid and ask rates for the world’s most frequently traded currencies: USD/EUR, USD/GBP, and USD/JPY over the sample period of 07-January-2014 to 30-September-2020. Using functional time-series models, we precisely forecast FX intraday volatilities and highlight their practical benefits through applications in intraday risk management.
@@
The intraday evolution of FX rate movements is conveniently cast as a functional data process in acknowledgement of the continuous underlying dynamic driving its behaviour. By observing FX intraday return curves, we find they are long-range dependent at the second moment and highly cross-correlated. We then methodologically innovate beyond prior functional GARCH-type frameworks proposed by Hörmann et al. (2013), Aue et al. (2017), and Cerovecki et al. (2019) by incorporating cross-dependence structure and long-range conditional heteroscedasticity features observed in the intraday FX data. Furthermore, we also contribute by extending traditional GARCH-style models to integrate bid-ask market microstructure information using the functional GARCH-X model. Note that our modelling framework can be generally applied to other financial markets for intraday volatility forecasting to provide risk management implications.
@@
As a result, we demonstrate the statistical superiority of our proposed models using the model confidence set and the Diebold-Mariano test. Taking into account the cross-currency dependence structure improves the accuracy of forecasting FX intraday conditional volatility. This conclusion is based on the MFPCA data-driven method that provides a superior basis for the dimension reduction process of the functional GARCH-type models, especially for the FGARCH(1,1) model. The success of the MFPCA basis may be driven by increased integration among economies and currency pairs, with each country’s external imbalances being a possible explanation for the observed cross-sectional variation (Della Corte et al., 2016). Moreover, explaining the long-range dependence is beneficial in forecasting intraday and inter-daily conditional volatility. The LFPCA basis used in FGARCH-type models moderately improves forecasting accuracy. More significantly, our results provide strong evidence to suggest that incorporating the bid-ask spread information through an FGARCH-X model enhances the predictability of conditional volatility of FX returns. The proposed models also show some outperformance in terms of inter-daily volatility forecasts compared with classic realised volatility-based models.
@@
We also demonstrate how one might economically benefit from our intraday volatility forecasting models. We calculate intraday VaR measures through intraday conditional volatility forecasts and further show how a VaR-corrected trading strategy might protect investors. Almost systematically across all currencies and forecasting models, we find superior returns, Sharpe ratios, and maximum drawdowns through the determination of the intraday VaR risk management signal, particularly with the MFPCA basis-based models consistently outperforming.
@@
For a potentially fruitful future research avenue, one might incorporate intraday options data sources into the functional GARCH-X framework, as option prices are often interpreted as the market’s expectation of future realised volatility. Exploring intraday volatility modelling and forecasting in other related financial markets will also be interesting. One could also extend the FGARCH models by applying a cube root transformation to the intraday return data (Noguchi et al., 2016), which would yield a more symmetric underlying distribution and potentially lead to more precise model estimation.
Acknowledgment
The authors would like to thank the Editor, an Associate editor and two reviewers for their instructive comments, which helped to improve both the substance and presentation of the paper. The second author is supported by an Australian Research Council Discovery Project (DP230102250). The authors are grateful for the comments from the participants at the Time Series and Forecasting Symposium held at the University of Sydney in 2022.

### v3

@@ 6 Conclusion
This paper seeks to enhance intraday FX market predictions by incorporating cross-asset dependence and long-range conditional heteroskedasticity features into return curve forecasting models. Our dataset includes the intraday 5-minute frequency spot price and bid and ask rates for the world’s most frequently traded currencies: USD/EUR, USD/GBP, and USD/JPY over the sample period of 07-January-2014 to 30-September-2020. Using FGARCH-type models, we precisely forecast FX intraday volatilities and highlight their practical benefits through applications in intraday risk management.
@@
The intraday evolution of FX rate movements is conveniently cast as a functional data process in acknowledgement of the continuous underlying dynamic driving its behaviour. By observing FX intraday return curves, we find that they are long-range dependent at the second moment and highly cross-correlated. We then methodologically innovate beyond the previous functional GARCH-type frameworks proposed by Aue et al., (2017) and Cerovecki et al., (2019) by incorporating the cross-dependence structure and the long-range conditional heteroskedasticity features observed in the intraday FX data. Furthermore, we also contribute by extending traditional GARCH-style models to integrate bid-ask market microstructure information using the functional GARCH-X model. Note that our modelling framework can be generally applied to other financial markets for intraday volatility forecasting to provide risk management implications.
@@
As a result, we demonstrate the statistical superiority of our proposed models using the model confidence set and the Diebold-Mariano test. Taking into account the cross-currency dependence structure improves the accuracy of forecasting FX intraday conditional volatility. This conclusion is based on the MFPCA data-driven method that provides a superior basis for the dimension reduction process of functional GARCH-type models, especially for the FGARCH(1,1) model. The success of the MFPCA basis may be driven by increased integration among economies and currency pairs, with each country’s external imbalances being a possible explanation for the observed cross-sectional variation (Corte et al.,, 2016). Moreover, explaining the long-range dependence is beneficial in forecasting intraday and inter-daily conditional volatility. The LFPCA basis used in FGARCH-type models moderately improves forecasting accuracy. More significantly, our results provide strong evidence to suggest that incorporating the bid-ask spread information through an FGARCH-X model enhances the predictability of conditional volatility of FX returns. The proposed models also show some outperformance in terms of inter-daily volatility forecasts compared with classic realised volatility-based models.
@@
We also demonstrate how one might economically benefit from our intraday volatility forecasting models. We calculate intraday VaR measures through intraday conditional volatility forecasts and further show how a VaR-corrected trading strategy might protect investors. Almost systematically across all currencies and forecasting models, we find superior returns, Sharpe ratios, and maximum drawdowns through the determination of the intraday VaR risk management signal, particularly with MFPCA basis-based models consistently outperforming.
@@
For a potentially fruitful future research avenue, one might incorporate intraday options data sources into the functional GARCH-X framework, as option prices are often interpreted as the market’s expectation of future realised volatility. Exploring intraday volatility modelling and forecasting in other financial assets with distinct characteristics represents an interesting direction for future research. One could also extend the FGARCH models by applying a cube root transformation to intraday return data, which would yield a more symmetric underlying distribution and potentially lead to more precise model estimation.